OpenStudy (anonymous):

The table shows the height of a plant as it grows. What equation in point-slope form gives the plant’s height at any time? Let y stand for the height of the plant in cm and let x stand for the time in months A. y – 15 = (x – 3) B. y – 15 = 5(x – 3) C. y – 3 = (x – 15) D. The relationship cannot be modeled.

OpenStudy (anonymous):

Time (months) Plant Height (cm) 3 15 5 25 7 35 9 45

OpenStudy (anonymous):

The table is posted right above.

OpenStudy (anonymous):

Pick any two lines and calculate the gradient. i will pick first 2 for it m = (25-15)/(5-3) m = 5 y-y1 = m(x-x1) y -15 = 5( x-3)

OpenStudy (anonymous):

The answer would be A

OpenStudy (anonymous):

Okay, I thought so but really wasnt sure. @AleshGames. Thank you!

OpenStudy (anonymous):

All Good :)